the Determine the area of the region y 1 + x² and y=xt3 region boanded by
Answer #1

9 =X+3 8 =X+3 To find the point A & B 9= 1+12 +х=х+3 * - 2 со --23 + - 2 со (1-2) +1 (-2) го G-2)(х+1) го d=1+12- 2 X= 2 - -​​​